# Dynamical de Sitter conjecture and quintessence

### v2

@@ I Introduction
On the other hand, the current accelerating cosmic expansion should be explained by a potential of the scalar field that lacks minimum or local minimum; If the potential remains consistently positive and the contribution from its kinetic term is not overly significant, one can still explain an accelerating universe consistent with current observational data. Such scalar fields are often referred to as quintessence Caldwell:1997ii ; Carroll:1998zi ; Zlatev:1998tr ; Chiba:1999ka ; Tsujikawa:2013fta . The simplest and most natural choice of the potential for quintessence models is given by an exponential function Barreiro:1999zs ; Raveri:2018ddi
@@
where $\tau$ is a scalar field and $\lambda$ is a constant slope. However, a puzzle issue has recently emerged with the quintessence models with the above scalar potential. The allowed region $\lambda<0.60$ of the potential slope $\lambda$ in the numerical solutions of the field equations in four-dimensional dynamical systems Andriot:2024jsh ; Rudelius:2022gbz ; Shiu:2023nph ; Copeland:1997et ; Ferreira:1997hj ; Bahamonde:2017ize ; Alestas:2024gxe ; Casas:2024oak suggested by observational facts of supernovae DES:2024jxu and BAO DESI:2024mwx is not consistent with the result $\lambda\geq\sqrt{2}$ constrained bythe de Sitter conjecture Obied:2018sgi . However, when considering quintessence, vanishing the time derivative of the scalar field is not necessarily a condition that must be satisfied. Here there is also a problem with quintessence as dark energy/late time accelerated expansion is that it is guaranteed to make $H_{0}$ tension. These issues are explored in Colgain:2019joh ; Banerjee:2020xcn ; Lee:2022cyh . They slowly generalized the statement from the potential (2), to generic potentials, and beyond quintessence to the Horndeski class of models.
There are some comments on the contents other than those mentioned above. For a multi-field scenario, the late-time attractor is necessarily the scaling solution Collinucci:2004iw such that the scale factor grows like a power of cosmological time Shiu:2023fhb ; Shiu:2024sbe . When the lower-dimensional dilaton is one of the rolling fields then the strong de Sitter conjecture is always fulfilled in the presence of just positive potentials Shiu:2023fhb ; Shiu:2023nph . There is also simple criteria even if we have negative potential terms Shiu:2023nph ; VanRiet:2023cca . One can also has some constraints about contracting universes with negative-definite potentials Shiu:2023yzt ; Uzawa:2018sal .
@@
II de Sitter conjecture
@@
II.1 Extension of Swampand criteria
Now we suppose a theory in $D$ dimensions and compactify it on a $(D-d)$-dimensional space with a metric $g_{mn}$ and a warp factor $\Omega$ as,
@@ II.1 Extension of Swampand criteria
where $g_{mn}$ denotes the metric of internal space, $x^{M}$ , $x^{\alpha}$  and $y^{m}$  are the coordinates of $D$-dimensional spacetime, $(d-1)$-dimensional space, $(D-d)$-dimensional internal space, respectively. We assume that the warp factor $\Omega$ and the internal space metric $g_{mn}$ depend only on the coordinates of the internal space $y^{m}$ and time $t$,.
@@
where $\bar{{\triangle}}$ indicates $(D-d)$-dimensional Laplace operator with respect to the metric $g_{mn}$. We will try to obtain a restriction on the values of the acceleration using the strong energy condition in $D$-dimensions. It gives
@@
Here, the tilde denotes that the overall modulus contribution was singled out. The relevant part of the $d$-dimensional effective action is
@@
$g_{mn}(t,y)$ $=$ $\exp\left[\sqrt{\frac{d-2}{(D-d)(D-2)}}\,\tau\right]\tilde{g}_{mn}(\Phi,y)\,.$ (16b)
@@
$\left[\frac{-d^{3}+d^{2}(D-1)+6D(d-2)+4}{4(D-2)(d-2)}\right]\dot{\tau}^{2}+\sqrt{\frac{D-d}{(D-2)(d-2)}}\kappa_{d}^{2}\partial_{\tau}V\,.$ (17)
@@
$-\frac{\partial_{\tau}V}{V}\geq 2\sqrt{\frac{D-2}{(D-d)(d-2)}}-\alpha\frac{\dot{\tau}^{2}}{V}\,,$ (18)
@@
$-\frac{\partial_{\tau}V}{V}\geq 2\sqrt{\frac{D-2}{(D-d)(d-2)}}=\lambda_{\rm SEC}\,.$ (21)
@@ III Application to Quintessence
with $V_{0}$ and $\lambda>0$ arbitrary constants. It is easy to verify that Eqs. (12c) and (13) are satisfied by
@@
The value of the term $\frac{\dot{\tau}^{2}}{V}$ for each total dimension $D$ is listed in table 1 , where we set $d=4$ and $\kappa_{4}^{2}=1$ and $\lambda$ takes the smallest positive value in each dimension. One can see that as $D$ decreases, the kinetic term of the scalar field $\frac{\dot{\tau}^{2}}{V}$ cannot be neglected for any value of $V_{0}$
Table 1: We show the contribution of $\dot{\tau}^{2}/V$ for $d=4$ , $\kappa_{4}^{2}=1$ , when $\lambda$ is the smallest positive value. As the number of total dimension $D$ decreases, the term $\dot{\tau}^{2}/V$ becomes large.
@@
In a setup of string theory ($D=10,d=4$), $\lambda$ should satisfy
@@
Figure 1: The allowed region of the parameter $\lambda$ for $d=4$ is depicted in the case of $D=11$ (a), $D=10$ (b), and $D=5$ (c). One can recognize that the lower bound of $\lambda$ increases as the number of total dimension $D$ decreases.
@@ IV Summary and discussions
The dS conjecture Obied:2018sgi cannot hold because of of the problem with the Higgs potential that we discovered in Denef:2018etk . This has not yet been resolved (see also Choi:2018rze ). Our new criteria also fall victim to the Higgs potential problem.
In our paper, the strong energy condition is assumed in the higher-dimensional theory. We assume that there is no Higgs potential as described by Denef:2018etk in the higher-dimensional theory. If there is a Higgs potential in original higher-dimensional theory, we have to consider the equation of motion of the higher-dimensional Higgs field in the process of deriving equation (18) in our paper. However, if the Higgs potential of Denef:2018etk appears ”effectively” in the four-dimensional effective theory, the new criteria will not solve the Higgs problem. Even if the de Sitter criteria have a kinetic term of scalar field, the Higgs potential has a much larger contribution than the quintessence potential. The strong energy condition imposed in the paper will be violated at least at the extreme value of the Higgs potential. Even in this case, it would be ideal to choose the ansatz of the metric well and consider a compactification in which the Higgs and quintessence potentials are described by Denef:2018etk . Furthermore, if we impose a strong energy condition in the four-dimensional effective theory, it will give a constraint on $\lambda$. Regarding the second derivative of the potential of a scalar field, which is treated in Ooguri:2018wrx , it cannot be obtained simply by combining field equations and energy conditions. The dS criteria involving the second derivative are discussed using a simpler derivation of the asymptotic form and point out a potential loophole Hebecker:2018vxz .
@@ Acknowledgements.
We thank Arthur Hebecker for valuable comments. The work of MN is supported in part by JSPS KAKENHI [Grants No. JP22H01221] and the WPI program “Sustainability with Knotted Chiral Meta Matter (SKCM2)” at Hiroshima University (KN and MN). The work of K. U. is supported by Grants-in-Aid from the Scientific Research Fund of the Japan Society for the Promotion of Science, under Contract No. 16K05364 and by the Grant “Fujyukai” from Iwanami Shoten, Publishers.

### v3

@@ I Introduction
On the other hand, the current accelerating cosmic expansion should be explained by a potential of the scalar field that lacks minimum or local minimum; If the potential remains consistently positive and the contribution from its kinetic term is not overly significant, one can still explain an accelerating universe consistent with current observational data. Such scalar fields are often referred to as quintessence Caldwell:1997ii ; Carroll:1998zi ; Zlatev:1998tr ; Chiba:1999ka ; Tsujikawa:2013fta .
The simplest and most natural choice of the potential for quintessence models is given by an exponential function Barreiro:1999zs ; Raveri:2018ddi
@@
where $\tau$ is a scalar field and $\lambda$ is a constant slope. However, a puzzle issue has recently emerged with the quintessence models with the above scalar potential. The allowed region $\lambda<0.60$ of the potential slope $\lambda$ in the numerical solutions of the field equations in four-dimensional dynamical systems Andriot:2024jsh ; Rudelius:2022gbz ; Shiu:2023nph ; Copeland:1997et ; Ferreira:1997hj ; Bahamonde:2017ize ; Alestas:2024gxe ; Casas:2024oak suggested by observational facts of supernovae DES:2024jxu and BAO DESI:2024mwx is not consistent with the result $\lambda\geq\sqrt{2}$ constrained by the trans-Planckian censorship conjecture and the strong de Sitter conjecture Bedroya:2019snp ; Rudelius:2021oaz . However, when considering quintessence, vanishing the time derivative of the scalar field is not necessarily a condition that must be satisfied. Here there is also a problem with quintessence as dark energy/late time accelerated expansion is that it is guaranteed to make $H_{0}$ tension. These issues are explored in Colgain:2019joh ; Banerjee:2020xcn ; Lee:2022cyh . They slowly generalized the statement from the potential (2), to generic potentials, and beyond quintessence to the Horndeski class of models.
@@
II de Sitter conjecture and Extension of Swampland criteria
@@
Now we suppose a theory in $D$ dimensions and compactify it on a (Dd)(D-d)-dimensional space with a metric $g_{mn}$ and a warp factor $\Omega$ as,
@@
where $g_{mn}$ denotes the metric of internal space, $x^{M}$ , $x^{\alpha}$  and $y^{m}$  are the coordinates of $D$-dimensional spacetime, (d1)(d-1)-dimensional space, (Dd)(D-d)-dimensional internal space, respectively. We assume that the warp factor $\Omega$ and the internal space metric $g_{mn}$ depend only on the coordinates of the internal space $y^{m}$ and time $t$,.
@@
where $\bar{{\triangle}}$ indicates (Dd)(D-d)-dimensional Laplace operator with respect to the metric $g_{mn}$. We will try to obtain a restriction on the values of the acceleration using the strong energy condition in $D$-dimensions. It gives
@@
Here, the tilde denotes that the overall modulus contribution was singled out. In what follows, we assume that the moduli $\Phi$ is fixed and focus only on the dynamics of $\tau$. The relevant part of the $d$-dimensional effective action is
@@
$g_{mn}(t,y)$ $=$ $\exp\left[2\sqrt{\frac{d-2}{(D-d)(D-2)}}\,\tau\right]\tilde{g}_{mn}(\Phi,y)\,$, (16b)
where note that we assume here that the moduli $\Phi$ other than $\tau$ are fixed.
@@
$\left[\frac{-d^{3}+d^{2}(D-1)+6D(d-2)+4}{4(D-2)(d-2)}\right]\dot{\tau}^{2}+\sqrt{\frac{D-d}{(D-2)(d-2)}}\kappa_{d}^{2}\partial_{\tau}V\,$, (17)
Here, we assume that the moduli $\Phi$ other than $\tau$ are fixed, and for the integral over the extra-dimensional space, we have used (4).
@@
$\frac{\partial_{\tau}V}{V}\geq 2\sqrt{\frac{D-2}{(D-d)(d-2)}}-\alpha\frac{\dot{\tau}^{2}}{V}\,$, (18)
@@
$\frac{\partial_{\tau}V}{V}\geq 2\sqrt{\frac{D-2}{(D-d)(d-2)}}=\lambda_{\rm SEC}\,.$ (21)
@@ III Application to Quintessence
with $V_{0}$ and $\lambda>0$ arbitrary constants. It is easy to verify that Eqs. (12c) and (13) are satisfied by  [^1] This is not the only possible solution which have the attractive fixed point at the late time. Since this is correct for $\varepsilon<d-1$ , we find an upper bound to $\lambda$ . Note that there is a kination solution, as well as solutions going from the latter to the former upon setting the potential (22).
@@
The value of the term $\frac{\dot{\tau}^{2}}{V}$ for each total dimension $D$ is listed in table 1 , where we set $d=4$ and $\kappa_{4}^{2}=1$ and $\lambda$ takes the smallest positive value in each dimension. One can see that as $D$ decreases, the kinetic term of the scalar field $\frac{\dot{\tau}^{2}}{V}$ cannot be neglected for any value of $V_{0}$ .
Table 1: We show the contribution of $\dot{\tau}^{2}/V$ for $d=4$ , $\kappa_{4}^{2}=1$ , when $\lambda$ is the smallest positive value. Here we present $\dot{\tau}^{2}/V$ with respect to smallest value of $\lambda$ because it takes minimum for $\lambda<2$ . As the number of total dimension $D$ decreases, the term $\dot{\tau}^{2}/V$ becomes large.
@@
Let us define the function $f(\lambda)$ ,
$f(\lambda)=\lambda^{3}-(2\alpha\,\kappa_{d}^{2}+b)\lambda^{2}-\frac{4(d-1)}{d-2}\kappa_{d}^{2}\lambda+\frac{4(d-1)}{d-2}b\,$, (32)
we will find the allowed region of $\lambda$ fulfilling $f(\lambda)\leq 0$ for each theory. In a setup of string theory ($D=10,d=4$), $\lambda$ should satisfy
@@
Figure 1: The allowed region of the parameter $\lambda$ for $d=4$ satisfying $f(\lambda)\leq 0$ is depicted in the case of $D=11$ (a), $D=10$ (b), and $D=5$ (c). One can recognize that the lower bound of $\lambda$ increases as the number of total dimension $D$ decreases.
@@
Previous treatments of the de Sitter conjecture have directly applied the results of Obied:2018sgi to observational cosmology without adequate consideration of the underlying assumptions. While this approach remains valid in regimes where scalar field dynamics are negligible $\dot{\phi}=0$ , substantial discrepancies arise when field dynamics play a significant role, as in quintessence scenarios. Our analysis demonstrates that proper inclusion of field kinetic terms yields parameter spaces consistent with four-dimensional observational cosmology. This approach represents a conceptually important advancement and constitutes a genuinely novel contribution, providing a rigorous analytical framework previously unavailable in the literature.
@@ IV Summary and discussions
Regarding the accelerated expansion solutions in string theory, one correctly notes the inherent difficulty in realizing the current accelerating expansion of universe under strong energy conditions without incorporating scalar field dynamics. However, our analysis demonstrates that when the scalar field dynamics are properly included rather than neglected, the prospects for achieving cosmological acceleration are significantly enhanced. Our central argument is that in constructing four-dimensional accelerating cosmologies, accounting for scenarios where the kinetic contributions of fields such as quintessence remain non-negligible allows us to identify parameter regimes consistent with observational cosmology - a result that was not attained in Obied:2018sgi .
The dS conjecture Obied:2018sgi cannot hold because of the problem with the Higgs potential that we discovered in Denef:2018etk . This has not yet been resolved (see also Choi:2018rze ; Han:2018yrk ). Our new criteria also fall victim to the Higgs potential problem.
In our paper, the strong energy condition is assumed in the higher-dimensional theory. We assume that there is no Higgs potential as described by Denef:2018etk in the higher-dimensional theory. If there is a Higgs potential in original higher-dimensional theory, we have to consider the equation of motion of the higher-dimensional Higgs field in the process of deriving equation (18) in our paper. However, if the Higgs potential of Denef:2018etk appears ”effectively” in the four-dimensional effective theory, the new criteria will not solve the Higgs problem. Even if the de Sitter criteria have a kinetic term of scalar field, the Higgs potential has a much larger contribution than the quintessence potential. The strong energy condition imposed in the paper will be violated at least at the extreme value of the Higgs potential. Even in this case, it would be ideal to choose the ansatz of the metric well and consider a compactification in which the Higgs and quintessence potentials are described by Denef:2018etk . Furthermore, if we impose a strong energy condition in the four-dimensional effective theory, it will give a constraint on $\lambda$.
@@ Acknowledgements.
We thank Arthur Hebecker for valuable comments. The work of MN is supported in part by JSPS KAKENHI [Grants No. JP22H01221 and JP23K22492] and the WPI program “Sustainability with Knotted Chiral Meta Matter (WPI-SKCM2)” at Hiroshima University. The work of K. U. is supported by Grants-in-Aid from the Scientific Research Fund of the Japan Society for the Promotion of Science, under Contract No. 16K05364 and by the Grant “Fujyukai” from Iwanami Shoten, Publishers.